Hello,
Please help!

equation 5x - 2y = 10 find the missing coordinate ( 6,a)
Thanks

Answer by stanbon(75887)

equation 5x - 2y = 10 find the missing coordinate ( 6,a)
Substitute to get:
5*6 -2a = 10
-2a = -20
a = 10
==========
Cheers,
Stan H.